# Quantifying Misalignment Between Agents: Towards a Sociotechnical Understanding of Alignment

### v2

@@ D.4 Overall Misalignment Across Problem Areas
$Pr(0|\mathcal{A},\mathcal{P}):=\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}\Big{(}Pr(a_{1},a_{2}\\ \text{ selected randomly from }\mathcal{A})\cdot c(g^{1}_{p},g^{2}_{p})\cdot\sqrt{w^{1}_{p}\cdot w^{2}_{p}}\Big{)}$
where $w^{i}_{p}$ is the weight agent $i$ assigns to problem area $p$.
@@
    $Pr(0|\mathcal{A},p):=Pr(a_{1},a_{2}\text{ selected randomly from }\mathcal{A})\\ \cdot c(g^{1}_{p},g^{2}_{p})$
@@
    Justification for geometric mean:
@@
    Justification for arithmetic mean:
    • It gives equal importance to each problem area in the overall score, before taking the agents’ importance weights into account.
@@
D.5 Misalignment Convergence
Theorem 5: As the number of agents $n$ approaches infinity, the misalignment score converges to a stable value independent of the number of problem areas, given that all problem areas have the same number of goals and conflict structure.
Proof: Let $\mathcal{G}$ be the number of goals in each problem area, and $c(i,j)$ be the conflict between goals $i$ and $j$.
  1. 1.
    As $n\to\infty$, by the law of large numbers, the proportion of agents with goal $i$ in any problem area converges to $\frac{1}{\mathcal{G}}$ for all $i$.
  2. 2.
    The misalignment score for a single problem area is:
    $M=\sum_{i=1}^{\mathcal{G}}\sum_{j=1}^{\mathcal{G}}\left(\frac{1}{\mathcal{G}}\cdot\frac{1}{\mathcal{G}}\cdot c(i,j)\right)$
  3. 3.
    This can be simplified to:
    $M=\frac{1}{\mathcal{G}^{2}}\sum_{i=1}^{\mathcal{G}}\sum_{j=1}^{\mathcal{G}}c(i,j)$
  4. 4.
    Since the conflict structure is the same for all problem areas, this value $M$ is identical for each problem area.
  5. 5.
    The overall misalignment across problem areas (Theorem 4) is the average of individual problem area misalignments:
    Overall $\text{Overall }M=\frac{1}{|\mathcal{P}|}\sum_{p=1}^{|\mathcal{P}|}M=M$
    where $|\mathcal{P}|$ is the number of problem areas.
  6. 6.
    Therefore, as $n\to\infty$, the overall misalignment converges to $M$, which is independent of the number of problem areas.
D.6 Goal Count and Misalignment Relationship
Theorem 6: For a problem area with $\hat{\mathcal{G}}$ non-zero goals, as the number of agents $n$ approaches infinity, the misalignment score approaches:
$\lim_{n\to\infty}\text{Misalignment}=\frac{\hat{\mathcal{G}}-1}{\hat{\mathcal{G}}}$ (3)
Proof: This follows directly from Theorem 3:
  1. 1.
    As $n\to\infty$, the distribution of agents across goals approaches uniformity.
  2. 2.
    In this limit, the probability of selecting two agents with different goals is $\frac{\hat{\mathcal{G}}-1}{\hat{\mathcal{G}}}$.
  3. 3.
    For mutually exclusive and maximally conflicting goals, this probability equals the misalignment score.
D.7 Maximum Misalignment with Uniform Distribution
Theorem 7: For a problem area with $\hat{\mathcal{G}}$ non-zero goals, the maximum misalignment occurs when the proportion of agents assigned to any single goal is $1/\hat{\mathcal{G}}$.
Proof: Let $p_{i}$ be the proportion of agents with goal $i$. The misalignment is:
$M=1-\sum_{i=1}^{\hat{\mathcal{G}}}p_{i}^{2}$ (4)
subject to the constraint $\sum_{i=1}^{\hat{\mathcal{G}}}p_{i}=1$.
Using the method of Lagrange multipliers:
  1. 1.
    Form the Lagrangian: $L=1-\sum_{i=1}^{\hat{\mathcal{G}}}p_{i}^{2}-\lambda(\sum_{i=1}^{\hat{\mathcal{G}}}p_{i}-1)$
  2. 2.
    Set partial derivatives to zero: $\frac{\partial L}{\partial p_{i}}=-2p_{i}-\lambda=0$ for all $i$
  3. 3.
    This implies all $p_{i}$ are equal. Given the constraint, $p_{i}=1/\hat{\mathcal{G}}$ for all $i$.
Therefore, the uniform distribution maximizes misalignment.
Appendix E Hardware and Software
All experiments were run on the following hardware:
@@ Appendix E Hardware and Software
    CPU: AMD Ryzen 9 5900HX with Radeon Graphics 3.30 GHz

### v3

@@ B.4 Overall Misalignment Across Problem Areas
$Pr(1|\mathcal{A},\mathcal{P}):=$ (13)
$\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}$ $\Big{(}Pr(a_{1},a_{2}\text{ selected randomly from }\mathcal{A}$
$a_{1}\neq a_{2})\cdot c(g^{1}_{p},g^{2}_{p})\cdot\sqrt{w^{1}_{p}\cdot w^{2}_{p}}\Big{)}$
where $w^{i}_{p}$ is the weight agent $i$ assigns to problem area $p\in\mathcal{P}$.
@@
    $Pr(1|\mathcal{A},p):=$ $\,\,Pr(a_{1},a_{2}\text{ selected randomly from }\mathcal{A},$
    $a_{1}\neq a_{2})\cdot c(g^{1}_{p},g^{2}_{p})$
@@
    We use the geometric mean because:
@@
    We use the arithmetic mean because:
    • It gives equal importance to each problem area in the overall score, such that the importance weights factored into the misalignment scores are what determines the overall importance of each problem area.
@@
Appendix C Applications of Misalignment Models
Current state-of-the-art in alignment methods used to develop frontier AI models lack consistency, theoretical backing, and sociotechnical grounding. For example, attempts to align foundation models to a vague notion of “humanity” have in practice led to models that mainly align to Western culture and values (Johnson et al. 2022). By defining and measuring misalignment in this more complex and nuanced way, we present a strong theoretical basis to better measure and track the success (or lack thereof) of alignment methods in the real world; avoid reductionist traps that assume that humans are always aligned with each other; and verify that they work as intended.
Intuitively, research and policy that intends to “align AI” will be more successful if it is conducted with a clear understanding of what the term “align” means, what success vs. partial success vs. failure looks like, and how to measure the above. Measuring alignment in an oversimplified manner may lead to mistaken assumptions that AI has been aligned where the reality is more complex – for example, AI systems aligned to one country’s cultural preferences could be misaligned with another population that has different preferences. As shown in §6.1: Shopping Recommender System, a misalignment analysis may reveal reward hacking scenarios as well as other concerns; our misalignment scores could potentially be used to improve system alignment by modifying certain aspects of the agentic setting. Our hope is to steer researchers away from reductive and oversimplified notions of alignment/misalignment and towards a richer sociotechnical understanding of alignment.
Other practical applications include designing new AI training methods for pluralistic alignment (e.g. applying multi-objective optimization methods to balance the preferences of multiple groups of annotators) and supplying policymakers with better levers for AI alignment governance (e.g. enabling the writing of policies that require new frontier AI models to achieve certain alignment scores on some set of problem areas and preference data).
Appendix D Hardware and Software
Simulating our model is, computationally, very cheap. All experiments were run on a commercial laptop with the following hardware:
@@ Appendix D Hardware and Software
    CPU: AMD64 Family 25 Model 80 Stepping 0, AuthenticAMD